# Post-detection inference for sequential changepoint localization

### v1

@@ 1 Introduction
meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$. We show how to construct such sets $\mathcal{C}$ for both types of algorithms (those with bounded ARL or PFA).
@@
Sequential (also known as the online or quickest) changepoint detection is a well-studied field, where the goal is to detect that a change has occurred as soon as possible, is crucial in applications requiring timely responses to distributional shifts, such as quality control, cybersecurity, medical diagnostics, and so on. The field traces back to early work on statistical process control. Shewhart, (1925) introduced the concept of a control chart, which computes a test statistic for each sample, with values within a threshold indicating the process remains in control. Page, (1954) proposed the CUSUM (cumulative sum) procedure, which computes the cumulative log-likelihood ratio under the assumption of two exactly known distributions. Another influential approach is the Shiryaev-Roberts (SR) procedure (Shiryaev,, 1963; Roberts,, 1966). These early methods, however, assume the pre- and post-change distributions to be fully specified. To relax the assumption of known post-change distributions, the Generalized Likelihood Ratio (GLR) procedure was introduced (Lorden,, 1971; Siegmund and Venkatraman,, 1995). Tartakovsky et al., (2014) provide comprehensive overviews of these classical methods, including CUSUM, SR, and GLR, and their numerous variants. Several nonparametric change-point detection algorithms have been developed, such as kernel-based methods (Harchaoui et al.,, 2008; Li et al.,, 2015) and e-detectors (Shin et al.,, 2023).
@@
Localizing the changepoint is another fundamental and crucial problem. As far as we are aware, the existing approaches to changepoint localization focus on offline or batch settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times. Hinkley, (1970) shows some asymptotic results about the maximum likelihood estimator (MLE) of the changepoint and also provides a confidence set based on likelihood ratio statistics. Yao, (1987) provides an approximation of the distribution of the MLE of the changepoint. Darkhovskh, (1976) presents a nonparametric changepoint estimator based on Mann-Whitney statistics. Carlstein, (1988) proposes a class of strongly consistent estimators for the changepoint in the nonparametric setting. Worsley, (1986) develops likelihood ratio-based confidence regions and tests for a changepoint for exponential family random variables. Dumbgen, (1991) constructs nonparametric confidence regions for changepoint by testing for each candidate time-point $t$ at level $\alpha$ using bootstrapping, achieving asymptotic coverage of $1-\alpha$, as the sample size $n\to\infty$. A more recent method by Jang and Walther, (2024) constructs confidence sets by evaluating a local test statistic on a triplet of time points and combining the results of the local tests with a weighted Bonferroni correction. To re-emphasize, all these methods are for the offline setting, while our main focus is on the sequential problem of constructing confidence sets after stopping, which has been largely overlooked in the literature. This gap is significant because in most real-world applications, data arrives sequentially, and making timely inferences about the changepoint immediately upon detecting a change is essential. Thus, developing methodologies for post-detection inference in a sequential framework is a significant step forward in changepoint analysis.
@@
The key idea behind constructing the confidence interval is simple: for each candidate $t\in\{1,\cdots,\tau\}$, we will test whether $t$ is the true changepoint and form the confidence set by inverting these tests. Our approach is conceptually similar to Dumbgen, (1991), it adapts those ideas to the sequential setting in a nontrivial manner. For example, we show that for change detection algorithms that have finite average run length (ARL), it is impossible to obtain finite length confidence intervals with marginal coverage guarantees. Further, for $1-\alpha$ conditional coverage, we show that we need to test the $t$-th hypothesis at level $\alpha\times\mathbb{P}_{\infty}(\tau\geq t)$ (a quantity we do not know apriori), whereas testing for each $t$ at level $\alpha$ works in the offline setting. We develop simulation based approaches to get around these hurdles.
@@
There are two settings for the changepoint problem: non-partitioned and partitioned. The non-partitioned changepoint problem assumes that the data-generating distribution changes from one unknown distribution $F_{0}\in\mathcal{P}$ to another unknown distribution $F_{1}\in\mathcal{P}$, without requiring any pre-specified partitioning of $\mathcal{P}$ into pre- and post-change distribution classes (eg: detecting a change in a Gaussian mean from any value to any other value). In contrast, the partitioned setting assumes that we know some partition of $\mathcal{P}$ into $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ such that $F_{0}$ and $F_{1}$ are known to lie in $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively. This formulation is useful when changes within $\mathcal{P}_{0}$ are not of interest, and we only seek to detect shifts from $\mathcal{P}_{0}$ to $\mathcal{P}_{1}$ (eg: detecting a change in Gaussian mean from negative to positive). Our method is designed for the partitioned setting. While extending it to the non-partitioned case is an interesting direction, it is left for future work.
Our contributions.
@@
  • For detection algorithms that control PFA, we propose Algorithm 1 for constructing a confidence set for the changepoint and we establish its unconditional and conditional coverage guarantees in Theorem 2.2 and Corollary 2.6, assuming both pre- and post-change distributions to be known.
@@
  • We demonstrate that the approach is not overly conservative (and thus practical) by examining the achieved coverage and size properties of our constructed confidence sets in extensive simulations.
We point out again that for obtaining the aforementioned guarantees, we just need to have black-box access to the detection algorithm $\mathcal{A}$, meaning that we do not need to know its inner workings (e.g., it can be a heuristic algorithm), and mainly need to be able to run that algorithm on any stream of data.
Paper outline.
The rest of the paper is organized as follows. In Section 2, we construct our confidence sets for the changepoint for known pre- and post-change and study its unconditional and conditional coverage probabilities. Section 3 extend this approach to handle composite post-change, assuming the pre-change distribution to be known. In Section 4, we finally extend this approach to handle composite pre- and post-change setting. Section 5.1 and Section 5.2 present a comprehensive set of simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach. This article is concluded in Section 6. Omitted proofs can be found in Appendix A. Some additional experimental results are provided in Appendix B and some heuristic algorithms are presented in Appendices C and D, which are computationally more efficient. A summary of comparisons of the heuristic and non-heuristic methods is provided in Appendix E, pointing at open avenues for further theoretical work.
@@ 2 Known pre- and post-change distributions
Notation.
For $t\in\mathbb{N}$, let $\mathbb{P}_{t}$ denote the joint distribution of the sequence of independent random variables $X_{1},X_{2},\dots$ when there is a changepoint at time t, meaning that $X_{1},\dots,X_{t-1}$ are from $F_{0}$ (with density $f_{0}$) and $X_{t},X_{t+1},\dots$ are from $F_{1}$ (with density $f_{1}$). As corner cases, let $\mathbb{P}_{1}$ denotes the joint distribution where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}$ and $\mathbb{P}_{\infty}$ denotes the case where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}$. Let $\mathbb{E}_{t}$ denote the expectation under $\mathbb{P}_{t}$, for $t\in\mathbb{N}\cup\{\infty\}$. Suppose we have a data sequence $\{X_{n}\}\sim\mathbb{P}_{T}$ for some unknown $T\in\mathbb{N}\cup\{\infty\}$ and some sequential change detection algorithm $\mathcal{A}$ that raises an alarm at the stopping time $\tau$, which is finite for the observed sequence.
In sequential change detection, there are typically two types of Type I error metrics. One is the average run length (ARL), which is defined as $\mathbb{E}_{\infty}[\tau]$, the expected time until a false alarm occurs under $T=\infty$ (i.e., no change). The other one is the probability of false alarm (PFA), that is, $\mathbb{P}_{\infty}(\tau<\infty)$. We will provide solutions for algorithms $\mathcal{A}$ that control either metric.
@@ 2.1 A first confidence set
Our main idea is to test whether there is a changepoint at $t$, for each $t\in\{1,\cdots,\tau\}$ and then construct the confidence set consisting of all time points $t$ for which we fail to reject the tests. Formally, we test the following null hypothesis at each candidate time point $t$:
@@
We consider some test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ such that for each $t\in\mathbb{N}$,
@@
1 Detect change at $\tau$ using $\mathcal{A}$
2 $\mathcal{C}=\varnothing$
@@ 11            $\mathcal{C}=\mathcal{C}\cup\{t\}$
We now provide more details about Subroutine $\mathcal{L}$.
@@ 2.2 Subtleties regarding Subroutine $\mathcal{L}$
The problem with $L=\infty$ is that the stopping time may never be reached for the $j$-th sampled sequence, i.e., we may have $\tau^{\prime}_{j}=\infty$. One can safely set $L=\infty$ if the $\mathcal{A}$ (i.e., its stopping time) satisfies the condition $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$, meaning that if there is a change, $\mathcal{A}$ stops almost surely. This is a very mild assumption because any reasonable change detector should always eventually raise an alarm if there is a change. Indeed, all algorithms with finite ARL satisfy the condition, and indeed typically so do algorithms that control PFA. To avoid making even this mild assumption, one can simply pick $L<\infty$.
If $L<\infty$ and the $j$th sequence stopped because the algorithm $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual, but if $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it). Therefore, for $L<\infty$,
@@
The rationale behind setting $M^{j}_{t,\infty}=-\infty$ when $\tau^{\prime}_{j}=\infty$ is the following. If for all $j=1,\cdots,B$, we have $\tau_{j}=\infty$, then we want to reject $H_{0,t}$, which is achieved by setting $M^{j}_{t,\infty}=-\infty$.
@@ 2.3 Prototypical choice of test statistic $M_{t}$
and the maximum likelihood estimator of $T$ (Hinkley,, 1970; Yao,, 1987) is given by
@@ 2.4 Unconditional coverage when $\mathcal{A}$ controls PFA
When $T\ll$ ARL, we typically have $\mathbb{\mathbb{P}}_{\infty}(\tau<T)\approx 0$ and so $\mathbb{P}_{T}(T\in\mathcal{C})\gtrapprox 1-\alpha$. The confidence set constructed using our method will approximately achieve the desired unconditional coverage level $1-\alpha,$ in situations where the probability of false detections before the true changepoint is negligible. However, $T$ is not known to us, and hence, we are unable to produce a suitable bound on the unconditional coverage for the detection algorithms that almost surely stop at a finite time, even when there is no change.
@@ 2.5 Unconditional coverage when $\mathcal{A}$ has finite ARL
If the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely), then it is impossible for any confidence set $\mathcal{C}^{*}$, which is a subset of $\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$. The intuitive reason is that if $\tau$ is random and finite almost surely, it is possibly a false detection, i.e., $\tau<T$, in which case the set $\{1,\cdots,\tau\}$ itself does not cover $T$. The next proposition shows that.
@@
If a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies an unconditional coverage guarantee $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$, for some $\alpha\in(0,1)$, then $\mathbb{P}_{\infty}(\tau=\infty)\geq 1-\alpha$ (thus, the PFA of the algorithm is at most $\alpha$ and thus the ARL of the algorithm is infinite). Said differently, for any algorithm with finite ARL, it is impossible to construct a valid confidence set that is only a subset of $\{1,\dots,\tau\}$.
The proposition says that if we can have a confidence set with high unconditional coverage, then the detection algorithm is such that it never stops with high probability if there is no change.
@@ 2.7 Modified confidence set when $\mathcal{A}$ has finite ARL (or $\mathcal{A}$ is a blackbox)
where $\hat{\alpha}_{t}=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. Algorithm 2 contains a structured overview of the method.
@@ 1 $\mathcal{C}=\varnothing$
7       Compute some test statistic $M_{t}$ (e.g., as defined in (2.9)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
@@
13      Compute $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$

### v2

@@ 1 Introduction
meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$. We show how to construct such sets $\mathcal{C}$ for both types of algorithms (bounded ARL or PFA).
@@
Sequential (also known as the online or quickest) changepoint detection is a well-studied field, where the goal is to detect that a change has occurred as soon as possible, is crucial in applications requiring timely responses to distributional shifts, such as quality control, cybersecurity, medical diagnostics, and so on. The field traces back to early work on statistical process control. Shewhart [19] introduced the concept of a control chart, which computes a test statistic for each sample, with values within a threshold indicating the process remains in control. Page [16] proposed the CUSUM (cumulative sum) procedure, which computes the cumulative log-likelihood ratio under the assumption of two exactly known distributions. Another influential approach is the Shiryaev-Roberts (SR) procedure [21, 18]. These early methods, however, assume the pre- and post-change distributions to be fully specified. To relax the assumption of known post-change distributions, the Generalized Likelihood Ratio (GLR) procedure was introduced [15, 22]. Tartakovsky et al. [23] provide comprehensive overviews of these classical methods, including CUSUM, SR, and GLR, and their numerous variants. Several nonparametric change-point detection algorithms have been developed, such as kernel-based methods [7, 14] and e-detectors [20].
@@
Localizing the changepoint is another fundamental and crucial problem. As far as we are aware, the existing approaches to changepoint localization focus on offline or batch settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times. Hinkley [8] shows some asymptotic results about the maximum likelihood estimator (MLE) of the changepoint and also provides a confidence set based on likelihood ratio statistics. Yao [26] provides an approximation of the distribution of the MLE of the changepoint. Darkhovskh [3] presents a nonparametric changepoint estimator based on Mann-Whitney statistics. Carlstein [1] proposes a class of strongly consistent estimators for the changepoint in the nonparametric setting. Worsley [25] develops likelihood ratio-based confidence regions and tests for a changepoint for exponential family random variables. Dümbgen [5] constructs nonparametric confidence regions for changepoint by testing for each candidate time-point $t$ at level $\alpha$ using bootstrapping, achieving asymptotic coverage of $1-\alpha$, as the sample size $n\to\infty$. A more recent method [10] constructs confidence sets by evaluating a local test statistic on a triplet of time points and combining the results of the local tests with a weighted Bonferroni correction. To re-emphasize, all these methods are for the offline setting, while our main focus is on the sequential problem of constructing confidence sets after stopping, which has been largely overlooked in the literature. This gap is significant because in most real-world applications, data arrives sequentially, and making timely inferences about the changepoint immediately upon detecting a change is essential. Thus, developing methodologies for post-detection inference in a sequential framework is a significant step forward in changepoint analysis.
@@
The key idea behind constructing the confidence interval is simple: for each candidate $t\in\{1,\cdots,\tau\}$, we will test whether $t$ is the true changepoint and form the confidence set by inverting these tests. Our approach is conceptually similar to [5], it adapts those ideas to the sequential setting in a nontrivial manner. For example, we show that for change detection algorithms that have finite average run length (ARL), it is impossible to obtain finite length confidence intervals with marginal coverage guarantees. Further, for $1-\alpha$ conditional coverage, we show that we need to test the $t$-th hypothesis at level $\alpha\times\mathbb{P}_{\infty}(\tau\geq t)$ (a quantity we do not know apriori), whereas testing for each $t$ at level $\alpha$ works in the offline setting. We develop simulation based approaches to get around these hurdles.
@@
There are two settings for the changepoint problem: non-partitioned and partitioned. The non-partitioned changepoint problem assumes that the data-generating distribution changes from one unknown distribution $F_{0}\in\mathcal{P}$ to another unknown distribution $F_{1}\in\mathcal{P}$, without requiring any pre-specified partitioning of $\mathcal{P}$ into pre- and post-change distribution classes (eg: detecting a change in a Gaussian mean from any value to any other value). In contrast, the partitioned setting assumes that we know some partition of $\mathcal{P}$ into $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ such that $F_{0}$ and $F_{1}$ are known to lie in $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively. This formulation is useful when changes within $\mathcal{P}_{0}$ are not of interest, and we only seek to detect shifts from $\mathcal{P}_{0}$ to $\mathcal{P}_{1}$ (eg: detecting a change in Gaussian mean from negative to positive). Our method is designed for the partitioned setting. Extensions to the non-partitioned case are left for future work.
1.1 Our contributions
@@
  • For detection algorithms that control PFA, we propose Algorithm 1 for constructing a confidence set for the changepoint and establish its unconditional and conditional coverage (Theorem 2.2, Corollary 2.6), assuming known pre- and post-change distributions.
@@
  • We demonstrate that the approach is not overly conservative (and thus practical) by examining the achieved coverage and size of constructed confidence sets in extensive simulations.
We point out again that for obtaining the aforementioned guarantees, we only need to have black-box access to the detection algorithm $\mathcal{A}$. It need not have formal guarantees, we do not need to know its inner workings, and just need to be able to run $\mathcal{A}$ on any stream of data.
1.2 Paper outline
In Section 2, we construct our confidence sets for the changepoint for known pre- and post-change and study its unconditional and conditional coverage probabilities. Section 3 extend this approach to handle composite post-change, assuming the pre-change distribution to be known. In Section 4, we finally extend this approach to handle composite pre- and post-change setting. Section 5.1 and Section 5.2 present a comprehensive set of simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach. This article is concluded in Section 6. Omitted proofs can be found in Appendix A. An additional illustrative example using the Poisson distribution is given in Appendix B. Appendix B. Some additional experimental results are provided in Appendix C and some heuristic algorithms are presented in Appendices D and E, which are computationally more efficient. A summary of comparisons of the heuristic and non-heuristic methods is provided in Appendix E, pointing at open avenues for further theoretical work.
@@ 2 Known pre- and post-change distributions
Notation. For $t\in\mathbb{N}$, let $\mathbb{P}_{t}$ denote the joint distribution of the sequence of independent random variables $X_{1},X_{2},\dots$ when there is a changepoint at time t, meaning that $X_{1},\dots,X_{t-1}$ are from $F_{0}$ (with density $f_{0}$) and $X_{t},X_{t+1},\dots$ are from $F_{1}$ (with density $f_{1}$). As corner cases, let $\mathbb{P}_{1}$ denotes the joint distribution where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}$ and $\mathbb{P}_{\infty}$ denotes the case where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}$. Let $\mathbb{E}_{t}$ denote the expectation under $\mathbb{P}_{t}$, for $t\in\mathbb{N}\cup\{\infty\}$. Suppose we have a data sequence $\{X_{n}\}\sim\mathbb{P}_{T}$ for some unknown $T\in\mathbb{N}\cup\{\infty\}$ and some sequential change detection algorithm $\mathcal{A}$ that raises an alarm at the stopping time $\tau$, which is finite for the observed sequence.
In sequential change detection, there are typically two types of Type I error metrics. One is the average run length (ARL), which is defined as $\mathbb{E}_{\infty}[\tau]$, the expected time until a false alarm occurs under $T=\infty$ (i.e., no change). The other one is the probability of false alarm (PFA), that is, $\mathbb{P}_{\infty}(\tau<\infty)$. We provide solutions for algorithms $\mathcal{A}$ that control either metric.
@@ 2.1 A first confidence set
We will test whether there is a changepoint at $t$, for each $t\in\{1,\cdots,\tau\}$ and then construct the confidence set consisting of all time points $t$ for which we fail to reject. Formally, we test the following null hypothesis at each candidate time $t$:
@@
Consider some test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ such that for each $t\in\mathbb{N}$,
@@
1 Detect change at $\tau$ using $\mathcal{A}$; Set $\mathcal{C}=\varnothing$
@@ 2.2 Subtleties regarding Subroutine $\mathcal{L}$
The problem with $L=\infty$ is that the stopping time may never be reached for the $j$-th sampled sequence, i.e., we may have $\tau^{\prime}_{j}=\infty$. One can set $L=\infty$ if $\mathcal{A}$ (i.e., its stopping time) satisfies the condition $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely). This is a very mild assumption because any reasonable change detector will always eventually raise an alarm if there is a change. Indeed, all algorithms with finite ARL satisfy the condition, and indeed typically so do algorithms that control PFA. To avoid making even this mild assumption, one can simply pick $L<\infty$.
If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual, but if $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it). Therefore, for $L<\infty$,
@@
The rationale behind setting $M^{j}_{t,\infty}=-\infty$ when $\tau^{\prime}_{j}=\infty$ is the following. If $\tau_{j}=\infty$, for all $j=1,\cdots,B$, then we want to reject $H_{0,t}$, which is achieved by setting $M^{j}_{t,\infty}=-\infty$.
@@ 2.3 Prototypical choice of test statistic $M_{t}$
and the maximum likelihood estimator of $T$ [8, 26] is given by
@@ 2.4 Unconditional coverage when $\mathcal{A}$ controls PFA
When $T\ll$ ARL, we typically have $\mathbb{\mathbb{P}}_{\infty}(\tau<T)\approx 0$ and so $\mathbb{P}_{T}(T\in\mathcal{C})\gtrsim 1-\alpha$. The confidence set constructed using our method will approximately achieve the desired unconditional coverage level $1-\alpha,$ in situations where the probability of false detections before the true changepoint is negligible. However, $T$ is not known to us, and hence, we are unable to produce a suitable bound on the unconditional coverage for the detection algorithms that almost surely stop at a finite time, even when there is no change.
@@ 2.5 Unconditional coverage when $\mathcal{A}$ has finite ARL
If the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely even if there is no change), then it is impossible for any confidence set $\mathcal{C}^{*}$, which is a subset of $\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$. The intuitive reason is that if $\tau$ is finite almost surely, it is possibly a false detection, i.e., $\tau<T$, in which case the set $\{1,\cdots,\tau\}$ itself does not cover $T$. The next proposition formalizes this fact.
@@
If a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies an unconditional coverage guarantee $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$, for some $\alpha\in(0,1)$, then $\mathbb{P}_{\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL of the algorithm is infinite). Said differently, for any algorithm with finite ARL, it is impossible to construct a valid confidence set that is only a subset of $\{1,\dots,\tau\}$.
The proposition says that if we have a confidence set with high unconditional coverage, then the detection algorithm never stops with high probability if there is no change.
@@ 2.7 Modified confidence set when $\mathcal{A}$ has finite ARL (or $\mathcal{A}$ is a blackbox)
where $\hat{\alpha}_{t}=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$. Algorithm 2 contains a structured overview of the method.
@@ 1 $\mathcal{C}=\varnothing$
7       Compute some test statistic $M_{t}$ (e.g. (2.9)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
@@
13      Compute $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$